OpenStudy (anonymous):

please help... Simplify the quantity of 2 x plus 10 all over the quantity of x squared plus 4 times x minus 5. the quantity of x plus 5 over the quantity of x minus 1 2 over the quantity of x minus 1 2 times the quantity of x plus 5 over the quantity of x minus 1 2 over x

OpenStudy (johnweldon1993):

\[\large \frac{2x + 10}{x^2 + 4x - 5}\] We can simplify the bottom as \[\large \frac{2x + 10}{(x + 5)(x - 1)}\] If we factor out a 2 out of the top...we have \[\large \frac{2(x + 5)}{(x + 5)(x - 1)}\] see anything that can cancel?

OpenStudy (anonymous):

the x and 5??

OpenStudy (johnweldon1993):

Right \[\large \frac{2\cancel{(x + 5)}}{\cancel{(x + 5)}(x - 1)}\] which leaves us with \[\large \frac{2}{x - 1}\]

OpenStudy (anonymous):

so thats the answer?

OpenStudy (johnweldon1993):

Indeed....but do you see how I got there?

OpenStudy (anonymous):

i dont see where you got to 5 when it was +10....

OpenStudy (johnweldon1993):

Okay...so just focusing on that part.. \[\large 2x + 10\] what do each term have in common? well, they are both divisable by 2...if we divie 2x by 2...we get x....and if we divide 10 by 2...we get 5 so we can write that as \[\large 2(x + 5)\]

OpenStudy (anonymous):

but then wouldnt it be just x+5?

OpenStudy (johnweldon1993):

Not quite...because then where would that 2 have gone? that's why we have it written outside the parenthesis...the 2 isnt gone...it's merely just "to the side" until we need it :)

OpenStudy (anonymous):

oh ok lol thanks....

OpenStudy (johnweldon1993):

Because we know that if we distribute the 2 back into the parenthesis....we will get back to the 2x + 10 like we had \[\large 2(x + 5) = 2x + 10\] If that helps at all lol

OpenStudy (anonymous):

oh ok